### v1

# Filtering for Creativity: Adaptive Prompting for Multilingual Riddle Generation in LLMs

Abstract
Multilingual riddle generation challenges large language models (LLMs) to balance cultural fluency with creative abstraction. Standard prompting strategies—zero-shot, few-shot, chain-of-thought—tend to reuse memorized riddles or perform shallow paraphrasing. We introduce Adaptive Originality Filtering (AOF), a prompting framework that filters redundant generations using cosine-based similarity rejection, while enforcing lexical novelty and cross-lingual fidelity. Evaluated across three LLMs and four language pairs, AOF-enhanced GPT-4o achieves 0.177 Self-BLEU and 0.915 Distinct-2 in Japanese, signaling improved lexical diversity and reduced redundancy compared to other prompting methods and language pairs. Our findings show that semantic rejection can guide culturally grounded, creative generation without task-specific fine-tuning.
1 Introduction
Large Language Models (LLMs) excel across many natural language processing tasks but often falter in creative generation within multilingual settings (Zhang (2025); Ismayilzada et al. (2024)). This limitation is especially apparent in bilingual riddle generation, where outputs are frequently rote and easily found online. To address this, we propose Adaptive Originality Filtering (AOF)—a prompting framework that promotes cultural specificity and lexical novelty without fine-tuning. AOF employs a semantic rejection mechanism to discard formulaic outputs and iteratively prompt the model until it produces a response that is both original and culturally coherent. In contrast to existing strategies like zero-shot, few-shot Brown et al. (2020a), and chain-of-thought prompting Wei et al. (2023), which often yield paraphrased common riddles, AOF actively steers generation toward novelty and cultural fidelity.
Figure 1: This diagram illustrates the end-to-end pipeline for generating and validating riddles using LLMs(GPT-4o, R1, LLaMA). The process begins with enforcing constraints on novelty and structure. Model outputs are then checked for semantic similarity using MiniLM, with a threshold score of << 0.75 determining acceptance. Outputs that fail are regenerated (Retry), while accepted samples undergo final evaluation.
We ask: Can prompting alone enable culturally aware, original bilingual generation in LLMs? To explore this, we apply AOF to three state-of-the-art LLMs (GPT-4o, LLaMA 3.1, DeepSeek) across four language pairs, benchmarking against four standard prompting baselines. We evaluate using metrics for lexical diversity (Distinct-n), redundancy (Self-BLEU), bilingual semantic alignment (Cross-Lingual BERTScore), and syntactic validity. Results show that AOF significantly improves both originality and cross-cultural alignment. By injecting an external filtering loop into prompting, AOF offers a lightweight, scalable path toward more creative and culturally attuned multilingual generation (see Figure 1).
2 Related Work
Most prior research emphasizes riddle solving rather than generation. Panagiotopoulos et al. (2024) and Heavey et al. (2024) developed multilingual inference models for riddle solving but did not extend to generative tasks. Xu et al. (2022) incorporated cultural embeddings for Chinese riddle comprehension, while Smith et al. (2022) evaluated NLG models on static benchmarks without multilingual or generative scope.
Multilingual representation efforts like MUSE Lample & Conneau (2019), LASER Chen & Avgustinova (2021), and XLM-R Conneau et al. (2019) support cross-lingual transfer but not figurative or culturally grounded generation. Sentence-level encoders Xia et al. (2019); Gritta & Iacobacci (2021) improve alignment but lack capacity for creative language. Dufter (2021) noted that such embeddings often collapse metaphorical nuance essential to cultural generation.
Creative Language Generation
Creative generation research has largely focused on ideation and open-domain tasks Ma et al. (2025). Cox et al. (2021; 2023) used structured prompts for idea diversity, while Huang et al. (2023) improved multilingual output via cross-lingual-thought prompting without filtering. Constraint-based methods like Yang et al. (2022) and Laverghetta Jr et al. (2024) use decoding control or heuristics to boost diversity.
Structural prompting has also gained traction: de Wynter et al. (2023) introduced recursive meta-prompts, Verma et al. (2024) identified abstraction limits in ReAct prompting, and Liu et al. (2024) applied paraphrasing under semantic constraints. Atmakuru et al. (2024) benchmarked narrative creativity under fixed templates, highlighting rejection-based filtering as a viable originality enhancer.
Prompting for Originality and Structure
Standard approaches prioritize structure over creativity. Few-shot prompting Brown et al. (2020a; b) and CoT Wei et al. (2023) improve reasoning but often replicate training artifacts. Iterative methods like Self-Refine Madaan et al. (2023), Tree-of-Thought Yao et al. (2023), and Reflexion Krishna et al. (2023) enhance fluency and factuality through retries but rarely enforce novelty.
Constraint-driven prompting offers stronger generative control. COLD decoding Mou et al. (2022), EditCoT Wang et al. (2024), and Sketch-of-Thought Aytes et al. (2025) use template-based completions, while DeLorean Liu et al. (2023) integrates symbolic constraints. These techniques target structure but overlook cultural fidelity. In contrast, our method enforces novelty during decoding via cosine-based filtering and supports cross-lingual generation of culturally resonant riddles.
3 Methodology
This section describes the components of our multilingual riddle generation framework, including our originality-enforcing prompting method, a centralized automatic evaluation metric, human annotation protocols, experimental design, and fine-tuning our model.
3.1 Adaptive Originality Filtering (AOF) Prompting
Traditional prompting strategies such as Chain-of-Thought (CoT) and Few-Shot improve logical coherence but frequently fall short in producing structurally novel and semantically diverse riddles. Prior work has shown that LLMs often reproduce internet riddles or paraphrase entries from BiRdQA Zhang & Wan (2022), defaulting to overused scaffolds such as “I have…” and “What am I?”. To address this, we propose Adaptive Originality Filtering (AOF), a prompting method designed to enforce semantic novelty, syntactic variation, and cross-lingual fidelity.
Semantic Filtering.
Each generated riddle $r_{\text{gen}}$ is compared against a reference set $\mathcal{D}=\{r_{i}\}_{i=1}^{N}$ using cosine similarity in embedding space. A candidate is accepted only if:
$\mathcal{S}(r_{\text{gen}},\mathcal{D})<\theta,\text{ with }\theta=0.75.$
The full formulation and reasoning for .75 appears in Appendix J.1.
Rejection Sampling Loop.
AOF employs rejection sampling: if a candidate fails the novelty filter, it is discarded and regenerated. This process repeats until a valid output is produced or a retry cap is reached. See Appendix J.2 for pseudocode.
Prompt Constraints.
Prompts used in AOF enforce specific closure formats, grammatical variation, and discourage common answers. Template rules are detailed in Appendix J.
3.2 Experimental Setup
We evaluate generation across multiple LLMs and prompting strategies, comparing structural, semantic, and fluency metrics in five languages.
Models Evaluated.
We benchmark GPT-4o, LLaMA 3.1, and DeepSeek Reasoning (R1), all queried under identical settings (temperature 0.7, token limit 3000).
Prompting Strategies.
We compare Zero-Shot, Few-Shot, Chain-of-Thought, Adversarial Wallace et al. (2019); Ribeiro et al. (2018), and AOF prompting. Full details are provided in Appendix K.
Dataset.
BiRdQA Zhang & Wan (2022) provides a bilingual corpus of 15,365 riddles in English and Chinese. We sample balanced subsets and align multilingual answers.
Metrics.
Automatic metrics include Self-BLEU (repetition), Distinct-2 (diversity), BERTScore (alignment). Human validation is included in Appendix A.
3.3 Fine-Tuning of the GPT-4o Model
Objective and Motivation
This fine-tuning aimed to improve GPT-4o-2024-08-06’s performance on riddle comprehension and generation across multiple languages. Riddles involve more than surface-level matching—they require metaphor understanding, logical contradiction, and creative misdirection. Our goal was not only to raise answer accuracy but also to instill structural reasoning abilities.
Methodological Overview
We framed the task as a supervised multi-class classification problem using the BiRdQA dataset. Riddles were presented in a multiple-choice format, and the model was fine-tuned using cross-entropy loss. For complete details on dataset preprocessing, training setup, and training set expansion, see Appendix I.
Multiple-Choice Framing Overview
Riddles were framed in a four-option multiple-choice format to encourage fine-grained discrimination between plausible distractors. This setup shaped the model’s reasoning strategies and generalization capabilities. A full analysis of framing effects is provided in Appendix I.5.
Prompting Strategies
We evaluated five prompting strategies on the fine-tuned model: Zero-Shot, Few-Shot, Chain-of-Thought (CoT), Adversarial, and Adaptive Originality Filtering (AOF). These were held consistent with the pretrained experiments. See Table 21 for full prompt templates.
Model Comparison Overview
We compared the fine-tuned GPT-4o to several pretrained baselines—GPT-4o (pretrained), LLaMA 3.1, and DeepSeek R1—using identical prompts and evaluation metrics. For detailed results and methodological discussion, see Appendix I.4.
4 AOF Pretrained Evaluations
We evaluate riddle generation across three pretrained models using five prompting selecting riddle one example per method. This section blends automatic metrics with qualitative observations to assess metaphor, creativity, and syntactic variation. Representative riddles are provided in Appendices DG. For detailed quantitative results—token length, Self-BLEU, Distinct-2, and syntactic validity—see Appendix B.
4.1 English
GPT-4o
achieves moderate repetition (Self-BLEU: 0.413) as shown in Table 5 and high lexical diversity (Distinct-2: 0.852), balancing structural cohesion with surface novelty. Compared to LLaMA 3.1 (0.471 / 0.727) and DeepSeek R1 (0.339 / 0.845), its generation reflects a middle ground: less repetitive than LLaMA, but more structurally conventional than R1. The riddle in Row 1 of Table 6 incorporates contrastive metaphor in a coherent frame, supporting findings that figurative ambiguity paired with consistent syntax enhances interpretability Lakoff & Johnson (1980); Shutova (2013). This reflects controlled creativity, where lexical risk is constrained by structural regularity Binsted (1996).
LLaMA 3.1
displays the strongest phrasal variation among the three models, as evidenced by a higher Distinct-2 (0.727), though with a moderate increase in repetition (Self-BLEU: 0.471). In Row 2, the riddle demonstrates rhythmic structure and concept layering, leveraging symmetry and internal metaphor to guide reader inference. This aligns with cognitive accounts linking memorable riddles to rhythmic and conceptual salience Koestler (1964). The AOF method appears to help LLaMA 3.1 decouple fixed templates from overused lexical forms, allowing for recompositional variety without semantic drift Fauconnier & Turner (2002).
DeepSeek R1
exhibits the lowest repetition (Self-BLEU: 0.339) and highest token diversity (Distinct-2: 0.845), outperforming both GPT-4o and LLaMA 3.1 in novelty. The output in Row 3 exemplifies conceptual inversion and abstract framing, using oppositional imagery to support lateral interpretation. This mirrors classic riddle mechanics involving duality and semantic misdirection Koestler (1964). While diversity of this magnitude sometimes correlates with fluency degradation Zhang et al. (2021), AOF appears to regulate R1’s output sufficiently to retain syntactic legibility Xu et al. (2018), enabling expressive reformulations without loss of coherence.
4.2 Japanese
GPT-4o
GPT-4o’s performance on metrics like self-BLEU and Distinct-n under AOF ranks around average relative to standard baselines—reflecting more on metric saturation than prompting inadequacy (Yao et al. (2025), Schmidtová et al. (2024)). AOF resolves key flaws in traditional prompts, such as the egocentric “I”-imagery in chain-of-thought and overfitting in few-shot examples. These gains are not fully captured by current metrics. As shown in Table 10, GPT-4o under AOF adopts a distinct structure: a punchy first sentence followed by a more elaborate second, enhancing narrative pacing and engagement.
LLaMA 3.1
While LLaMA 3.1 shows limited improvement on automated metrics using AOF, this reflects model expressiveness more than prompt quality. AOF mitigates egocentric phrasing and rote repetition seen in baseline prompts, yielding subtler gains beyond surface metrics. For instance, Table 10 features a clever use of the homophone 「つる」 to suggest both decorative twine and the crane (鶴)—symbols deeply embedded in Japanese culture and Shinto ritual, such as しめ縄 (shimenawa) An (2023).
DeepSeek R1
DeepSeek R1’s middling scores on self-BLEU and Distinct-n under AOF say more about its stylistic tendencies than the prompt’s design Li et al. (2024). AOF corrects core issues in baseline methods—overuse of first-person voice in chain-of-thought and example mimicry in few-shot—enabling richer creativity not fully reflected in metrics. Table 10 presents a vivid example: a fish’s mouth described as a “quiet tree” where birds sing, merging the surreal and natural in a poetically disorienting twist.
4.3 Arabic
GPT-4o
GPT-4o shows moderate repetition (Self-BLEU: 0.497) as shown in Table 5 and good lexical variety (Distinct-2: 0.780) with Adaptive Originality Filtering (AOF), clearly performing better than common methods like few-shot, zero-shot, chain-of-thought, and adversarial prompts. Unlike chain-of-thought prompts, which tend to produce straightforward, predictable metaphors, AOF helps GPT-4o create riddles with imaginative and abstract images—such as something that’s present but unseen—as illustrated in (Figure 2, Row 1). This approach fits naturally with traditional Arabic riddles, known for their symbolic and reflective style Al-Khatib (1988).
LLaMA 3.1
LLaMA 3.1 strikes an effective balance between repetition (Self-BLEU: 0.374) and creativity (Distinct-2: 0.927) through AOF, addressing issues often found in chain-of-thought and adversarial prompts, which frequently yield predictable or overly vague outputs. Its riddles are relatable and culturally resonant, using clear metaphors drawn from everyday life, like ”a strong wind” that can’t enter a house, as shown in (Figure 2, Row 2). This connects directly to familiar poetic traditions in Arabic, avoiding common pitfalls like repetitive phrasing or loss of meaning Al-Jahiz (869).
DeepSeek R1
DeepSeek R1, while somewhat repetitive (Self-BLEU: 0.585), achieves notable depth in metaphorical expression (Distinct-2: 0.583) under AOF. This method effectively tackles problems seen in zero-shot, few-shot, and adversarial prompting, such as repetitive or simplistic metaphors. For example, DeepSeek R1 creatively portrays a rooftop as an eye ”fed by the city,” as seen in (Figure 2, Row 3), mixing urban imagery with striking visual symbolism. This clever blending of abstract ideas and real-world images strongly aligns with Arabic poetry, known for its layers of meaning and subtle metaphors Al-Marzouki (2012). By encouraging culturally rich riddles, AOF clearly boosts the originality and depth of DeepSeek R1’s outputs compared to simpler prompting strategies Xu et al. (2018).
4.4 French
GPT-4o
GPT-4o’s pretrained riddles are grammatically solid and easy to understand, but they often feel like lifted translations of English puzzles. For example, it offers the riddle in Row 1 of Table 17, a fluent but familiar “cloud” trope. The phrasing remains straightforward, with minimal use of inversion or enjambment that one might expect in classic French énigmes Chan (1996). Even when prompted for more creativity, GPT-4o tends to default back to elemental imagery—wind, water, shadows—rather than exploring urban or abstract concepts. Answerability is never in doubt, but the surface novelty and cultural resonance remain modest. This is reflected in a Self-BLEU of 0.413 and Distinct-2 of 0.852 as shown in Table 5, suggesting moderate repetition and relatively strong lexical variety.
DeepSeek R1
DeepSeek R1 generates concise and structurally consistent riddles, yet it leans heavily on classic “river” or “echo” formats. A typical example, shown in Row 2, feels like a direct adaptation of childhood puzzles, with little lexical or rhythmic innovation Meulemans (2005). Attempts to push deeper—such as referencing time, memory, or abstract concepts—often collapse back into familiar patterns. Even in adversarial or chain-of-thought modes, R1 seldom ventures beyond these safe metaphors. The result is always coherent but predictable, with cultural idioms and advanced metaphorical shifts under-utilized. This pattern yields a low Self-BLEU of 0.339 and a high Distinct-2 of 0.845, indicating minimal paraphrastic overlap and strong surface novelty.
LLaMA 3.1
LLaMA 3.1 shows the greatest stylistic range—some riddles stumble through literal phrasing, while others introduce intriguing wordplay. Row 3 plays with poetic misdirection (“danse” / “ris”), while Row 4 experiments with digital metaphor (“curseur”), reflecting a willingness to stretch genre boundaries. These flashes of originality are countered by examples that revert to stilted syntax or literal templates. Overall, LLaMA’s pretrained outputs blend moments of genuine creativity with occasional lapses into generic phrasing, suggesting strong potential but inconsistent performance. The variation is mirrored in its Self-BLEU score of 0.471 and Distinct-2 of 0.727, balancing moderate repetition with decent surface diversity Veale (2011).
4.5 Chinese
GPT-4o
GPT-4o’s pretrained Chinese riddles are grammatically correct and logically coherent, but they often translate English metaphors without adapting to the script-specific strategies typical of traditional 灯谜. As shown in Row 1 of Table 13, the imagery is literal and binary, missing multi-layered allusions like radical-based clues or idiomatic rhythm Chan (1996); Sun (2006). When prompted for variation, GPT-4o maintains syntactic fluency but rarely ventures into lexical innovation. With a Self-BLEU of 0.280 and Distinct-2 of 0.869 as in shown in Table 5, it balances mild repetition with surface diversity but lacks deeper cultural anchoring.
DeepSeek R1
DeepSeek R1 outputs elegant, fluent couplets with strong adherence to classical poetic symmetry, as shown in Row 2. While it maintains strong rhythm and antithesis, its metaphors remain shallow—favoring form over lexical novelty. Even when attempting radical-based clues (Row 3), the effort often feels literal rather than layered. Despite this, R1 exhibits high diversity (Distinct-2: 0.674) and moderate repetition (Self-BLEU: 0.433), suggesting a capacity for expressive phrasing that falls short of fully exploiting Chinese script-level wordplay Xu et al. (2018).
LLaMA 3.1
LLaMA 3.1 demonstrates the richest cultural range. Row 4 blends visual and semantic metaphor in a style reminiscent of folk riddles, while Row 5 shows explicit use of radical-based structure. This script-aware design reflects deeper integration with Chinese morphological conventions Li (2008). Although some outputs revert to generic “无…能…” templates or awkward logic, LLaMA’s outputs remain stylistically diverse (Distinct-2: 0.776) and moderately novel (Self-BLEU: 0.428). This suggests that the AOF prompt enables meaningful variation while retaining cultural fidelity Fauconnier & Turner (2002).
Language Pair Prompting Method Fine-Tuned GPT-4o (Self-BLEU / Distinct-2)
English–Arabic Few-Shot 0.233 / 0.826
AOF (Ours) 0.260 / 0.893
Zero-Shot 0.391 / 0.752
Chain-of-Thought 0.326 / 0.831
Adversarial 0.320 / 0.810
English–Chinese AOF (Ours) 0.163 / 0.934
Zero-Shot 0.315 / 0.831
Few-Shot 0.349 / 0.787
Chain-of-Thought 0.305 / 0.828
Adversarial 0.400 / 0.757
English–Japanese AOF (Ours) 0.177 / 0.915
Zero-Shot 0.431 / 0.752
Few-Shot 0.326 / 0.778
Chain-of-Thought 0.386 / 0.796
Adversarial 0.327 / 0.748
English–French Chain-of-Thought 0.256 / 0.892
AOF (Ours) 0.273 / 0.856
Zero-Shot 0.289 / 0.867
Few-Shot 0.323 / 0.835
Adversarial 0.359 / 0.793
Table 1: Prompting performance (Self-BLEU / Distinct-2) for the fine-tuned GPT-4o model across language pairs. One bolded cell per language pair shows the best combined performance (lowest Self-BLEU and highest Distinct-2).
5 Fine-Tuned vs. Pretrained Riddle Generation
We compare GPT-4o before and after fine-tuning across five prompting strategies. Quantitative metrics—token length, Self-BLEU, and Distinct-2—are complemented by qualitative analysis of metaphorical framing, structural variation, and bilingual phrasing. Representative pairs are shown in Appendices DG.
5.1 English
[…]
5.2 Japanese
[…]
5.3 Chinese
[…]
5.4 Arabic
[…]
5.5 French
[…]
6 Fine-Tuned Riddle Comparison to Real World
We evaluate fine-tuned riddle generations across five prompting strategies by comparing them to their closest literary or folkloric counterparts. This comparison focuses on structural fidelity, metaphorical depth, and cultural alignment. Language-specific analyses are provided in Appendix A, which details how fine-tuned outputs differ from real-world riddles in phrasing, abstraction, and stylistic form.
7 Conclusion
This paper introduced Adaptive Originality Filtering (AOF), a prompting framework designed to overcome originality limits in bilingual riddle generation. Evaluated across four language pairs, five prompting paradigms, and three LLMs, AOF consistently outperformed baselines in lexical diversity (Distinct-n), semantic novelty (Self-BLEU), and cross-lingual alignment. Fine-tuned GPT-4o with AOF produced the highest-quality outputs, outperforming both its pretrained version and open-weight models. Human evaluations confirmed that AOF-generated riddles were more metaphorically rich, structurally coherent, and culturally resonant—especially in French, Japanese, Chinese, and Arabic. These findings demonstrate that filtering-based prompting, particularly when combined with fine-tuning, provides a robust path toward semantically novel and linguistically sound multilingual generation. AOF marks a practical advance in prompt design for creative, cross-lingual NLP tasks.
Limitations
Dataset Scope
We restrict our study to the BiRdQA dataset, covering 6,614 English and 8,751 Chinese riddles with multiple-choice answers. While genre-diverse, its figurative focus narrows generalizability to broader creative tasks (e.g., storytelling, allegory). Our multilingual evaluation spans five languages (EN–ZH–AR–JA–FR), excluding lower-resource or morphologically rich ones like Finnish or Swahili.
Prompting and Sampling
We fix sampling parameters (e.g., temperature, token limit) across all models to ensure comparability, potentially masking prompt–parameter effects. AOF’s MiniLM-based filtering emphasizes surface-level novelty, risking semantic redundancy undetected by cosine distance—particularly in non-English outputs.
Fine-Tuning Setup
Our fine-tuning on GPT-4o uses BiRdQA’s multiple-choice framing, enhancing structural fluency but biasing toward riddle types favoring clarity over ambiguity. Although stylistic gains are evident in token length and Self-BLEU, we omit extrinsic human metrics like puzzle difficulty or solver accuracy.
Evaluation Constraints
Human evaluations cover only five languages due to cost, limiting cultural generalizability. Syntactic validity checks apply solely to AOF generations, leaving other strategies unchecked structurally. Lastly, while AOF excels in novelty-structure balance, deeper human validation (e.g., creativity ratings, solve rates) remains future work.
Ethics Statement
Language Equity and Cultural Representation
This work evaluates riddle generation across five languages—English, Chinese, Japanese, Arabic, and French—selected for linguistic diversity and resource availability. However, our dataset and prompts reflect biases inherent to internet-derived corpora and may not fully capture the cultural or idiomatic richness of underrepresented communities. For example, certain metaphorical forms or idiomatic patterns may be overrepresented in English and underdeveloped in Arabic or Japanese, despite efforts to balance qualitative evaluation across languages.
Creative Attribution and AI Authorship
Riddles generated in this study may closely resemble publicly available riddles from folklore, educational resources, or crowdsourced riddle repositories. Although our Adaptive Originality Filtering (AOF) method actively screens for semantic overlap, we acknowledge that latent training data exposures or structural mimicry may inadvertently echo human-created content. As such, we recommend that users and downstream applications of our system avoid commercial use or publication of model outputs without further originality assessment or attribution checks.
Data Privacy and Responsible Fine-Tuning
The BiRdQA dataset used for training and evaluation contains no personally identifiable information (PII). All riddles are anonymized and synthetically framed as general-knowledge metaphors. Our fine-tuning process adheres to OpenAI’s guidelines for API use and operates within approved token limits and safety thresholds. No user data or private content was used in training, generation, or evaluation.
Misuse Risks and Interpretability
While riddle generation appears benign, the creative ambiguity inherent in figurative tasks can be exploited for misinformation or manipulation, especially in politically or culturally sensitive contexts. Models capable of abstract generation may also fabricate plausible but misleading information. We caution against deploying this system in high-stakes educational, psychological, or legal settings without adequate interpretability safeguards or human oversight.
Acknowledgements
We thank Bakr Bouhaya for his thoughtful analysis of Arabic riddles, which provided valuable insights and cultural grounding for our study.
Appendix A Appendix: Fine-Tuned AOF Riddle Comparison to Real World
A.1 English
As shown in Table 8, Row 1, the fine‑tuned riddle reimagines the original with more abstract and layered associations. Rather than relying on negated literalism, it introduces concepts like memory and time using metaphorical compression and cross-sensory cues. This approach reflects principles of conceptual integration theory, where blending disparate domains enhances figurative depth Fauconnier & Turner (2002). In contrast, the real-world version is more direct, using structural opposition to achieve its effect Gentner (1983).Row 2 presents another clear shift in stylistic strategy. The real-world riddle uses static reversal—a common riddle trope—while the fine-tuned variant introduces paradox and disappearance as metaphors for guidance. This relies on spatial embodiment, a known technique in metaphor production Lakoff & Johnson (1980), and adds ambiguity through indirectness. Such compositional layering suggests that AOF prompts can foster more interpretively open and stylistically creative riddles than conventional examples.
A.2 Japanese
The riddles in AOF are guided towards direct metaphors with complex, creative, and unique word choice and sentence structure, while having creative answers like memory and beehive in Table 12 Teng & Xu (2023). These generations surpass past riddle generations flaws like lack of originality in sentence structure, just changing the pronouns or verbs to make it more creative, and etc. These riddles contrast with traditional Japanese riddles which rely on phonetic ambiguity and cultural nuance like in Table 12 where the first row features how phonetically similar words feature different meanings and the riddle in the second row yields different ways of reading through phonetically similar readingsAn (2023).
A.3 Chinese
Fine-tuned AOF riddles in Chinese often leverage character structure through radical-based puns and vivid imagery. For instance, the coral riddle in Table 15 blends “sea” imagery with radical hints (海底藏森林…) to guide the solver—a strategy supported by prior work on character-pun alignments in riddle composition Tan et al. (2016). By contrast, traditional 灯谜 (e.g., “口袋里有个圆…” for “月亮”) rely on simple perceptual clues and tonal balance Wei & Lee (2021). This comparison suggests that our approach enhances cultural depth by embedding multi‑layered orthographic play into poetic metaphors while preserving reader accessibility.
A.4 Arabic
(Figure 4, Row 5) AOF stands out for its fresh language and metaphorical clarity. One riddle—”Something that’s full when it eats, and thirsty when it drinks”—relies on a simple yet clever contradiction that invites reflection. It draws on the tradition of using everyday logic to confuse and amuse, evoking the style of oral riddles that play with basic physical experiences. The second riddle—”I light up the night and disappear by day, visible yet unseen… What am I?”—is more poetic, using contrast and imagery to express something elusive and symbolic. It captures the feel of classical Arabic alghāz not through root-based punning but through layered metaphor and rhythm. Together, these examples show how AOF preserves the spirit of traditional riddling through modern, metaphor-rich language Antar (2023); Bhatt & Kuka (2025); Liu et al. (2022).
A.5 French
Fine-tuned AOF riddles in French lean into unexpected domain shifts and internal echo. The AOF example repurposes the concept of a “typo” as a buzzing bee, combining internal rhyme (“jardin/des mots”, “bourdonnant/lettres”) and metaphorical layering, driving semantic playfulness and rhythmic balance (Table 19, Row 1). Internal rhyme notably enhances poetic cohesion and cognitive engagement Encyclopædia Britannica (2025). In contrast, canonical French énigmes tend toward binary negation and elemental imagery (Table 19, Row 2). For instance, “Je vole sans ailes, je pleure sans yeux…” relies on simple antithesis without cross-domain metaphorical transfer. The AOF variant’s richer conceptual mapping aligns with findings that cross-domain metaphor and internal structure boost interpretability and novelty in poetic forms Lakoff & Johnson (1999); Encyclopædia Britannica (2025).
Appendix B Appendix: Additional Results Tables
B.1 Average Token Length Across Pretrained Models
Language Pair Prompting Method GPT-4o LLaMA 3.1 DeepSeek R1
English–Arabic Chain-of-Thought 910 1613 1085
Zero-Shot 1112 1519 2005
Few-Shot 1921 2050 3144
Adversarial 938 2202 1826
AOF (Ours) 1548 1157 2138
English–Chinese Zero-Shot 702 731 719
Few-Shot 2030 2097 2351
Chain-of-Thought 942 1389 1205
Adversarial 916 950 1126
AOF (Ours) 1275 1663 1535
English–Japanese Zero-Shot 1099 1127 1115
Few-Shot 1922 1941 2330
Chain-of-Thought 1169 1099 1802
Adversarial 1101 894 1128
AOF (Ours) 1185 1230 1273
English–French Adversarial 787 1128 1413
Zero-Shot 1163 1183 1613
Few-Shot 2061 2982 2565
Chain-of-Thought 940 1631 1236
AOF (Ours) 1166 1517 1982
Table 2: Average token lengths for each model and prompting method across language pairs. Bold = shortest average length per pair.
B.2 Average Token Lengths Across Languages
Language Pair Prompting Method Fine-Tuned GPT-4o (Avg. Token Length)
English–Arabic AOF (Ours) 1129
Zero-Shot 799
Few-Shot 1999
Chain-of-Thought 730
Adversarial 737
English–Chinese AOF (Ours) 1034
Zero-Shot 898
Few-Shot 2150
Chain-of-Thought 860
Adversarial 785
English–Japanese AOF (Ours) 894
Zero-Shot 894
Few-Shot 2088
Chain-of-Thought 753
Adversarial 844
English–French AOF (Ours) 1076
Zero-Shot 943
Few-Shot 2005
Chain-of-Thought 733
Adversarial 716
Table 3: Average token lengths for fine-tuned GPT-4o. Bold = shortest per pair.
B.3 Cross-Lingual Evaluation of Syntactic Validity
Language Model Total Riddles Valid Structures Validity (%)
English (EN) GPT-4o-fine-tune 10 10 100.0%
Chinese (ZH) GPT-4o-fine-tune 10 10 100.0%
Japanese (JA) GPT-4o-fine-tune 10 10 100.0%
Arabic (AR) GPT-4o-fine-tune 10 10 100.0%
French (FR) GPT-4o-fine-tune 10 10 100.0%
Table 4: Cross-lingual evaluation of syntactic validity of GPT-4o AOF generations.
B.4 Average self-BLEU and Distinct-n Pretrained Metrics
Language Pair Prompting Method GPT-4o LLaMA 3.1 DeepSeek R1
English–Arabic AOF (Ours) 0.497 / 0.780 0.374 / 0.927 0.585 / 0.583
Zero-Shot 0.272 / 0.975 0.432 / 0.746 0.627 / 0.543
Few-Shot 0.272 / 0.880 0.432 / 0.746 0.627 / 0.543
Chain-of-Thought 0.375 / 0.756 0.575 / 0.643 0.330 / 0.793
Adversarial 0.330 / 0.798 0.342 / 0.727 0.672 / 0.507
English–Chinese AOF (Ours) 0.280 / 0.869 0.428 / 0.776 0.433 / 0.674
Zero-Shot 0.335 / 0.739 0.482 / 0.649 0.320 / 0.854
Few-Shot 0.640 / 0.420 0.660 / 0.440 0.650 / 0.450
Chain-of-Thought 0.363 / 0.777 0.403 / 0.815 0.430 / 0.767
Adversarial 0.363 / 0.820 0.593 / 0.570 0.466 / 0.735
English–Japanese AOF (Ours) 0.483 / 0.697 0.516 / 0.640 0.560 / 0.690
Zero-Shot 0.364 / 0.833 0.430 / 0.871 0.514 / 0.757
Few-Shot 0.280 / 0.844 0.587 / 0.605 0.402 / 0.715
Chain-of-Thought 0.532 / 0.697 0.447 / 0.753 0.500 / 0.630
Adversarial 0.334 / 0.794 0.599 / 0.586 0.405 / 0.741
English–French AOF (Ours) 0.413 / 0.852 0.471 / 0.727 0.339 / 0.845
Zero-Shot 0.451 / 0.833 0.476 / 0.715 0.520 / 0.849
Few-Shot 0.371 / 0.814 0.480 / 0.665 0.670 / 0.535
Chain-of-Thought 0.444 / 0.733 0.455 / 0.750 0.359 / 0.768
Adversarial 0.358 / 0.806 0.485 / 0.614 0.461 / 0.673
Table 5: Prompting performance (Self-BLEU / Distinct-2). Bold = best combined (low Self-BLEU + high Distinct-2).
Appendix C Appendix: English Riddle Examples
C.1 English Pretrained Riddle Generations
Table 6: Representative English riddles generated under AOF prompting across pretrained models.
Model Riddle (English) Answer
GPT-4o It waits behind every choice, seen only once it’s gone. It changes nothing, yet weighs more than stone. Regret
LLaMA 3.1 I do not shine, but I am light. I cannot burn, yet I spark insight. I have no tongue, yet I speak in waves. Idea
DeepSeek R1 I echo where silence should rest. I fill the void with imagined guests. I’m absent, yet I dwell in minds. Memory
C.2 English Comparison of Fine-Tuned Riddle Generations to Pretrained Counterparts
Table 7: English Example Riddles for Pre-trained vs. Fine-Tuned Generations
Prompting Method Pre-trained Example Riddle Fine-Tuned Example Riddle
Zero-Shot I have keys but open no locks; I have space but no room. You enter numbers, letters, and more. What am I? I run without legs, whisper without a mouth. Who am I?
Few-Shot I’m full of holes, yet I hold water. What am I? I drift on unseen roads, carrying rain-songs in my wake. What am I?
Chain-of-Thought I have cities, but no houses; forests, but no trees; rivers, but no water. What am I? Kingdoms without subjects, roads without dust; I exist only in paper trust.
AOF (Ours) What is so fragile that saying its name breaks it? Softly spoken yet never heard, I am the quietest word.
Adversarial I fly without wings, I cry without eyes. Wherever I go, darkness flies. What am I? I erase mountains grain by grain, yet thirst is a stranger to me. What am I?
C.3 English Fine-Tuned Riddles and Their Real-World Counterparts
Table 8: English Riddle Comparison: AOF Fine-Tuned vs. Real-World
Row Real-World Riddle AOF Fine-Tuned Riddle
1 I have hands but cannot clap. What am I? I carry time but never age. I never forget, but I cannot speak. What am I?
2 I guide people across the land, but I never move. What am I? I lead with no voice, move without steps, and vanish when sought. What am I?
C.4 English Fine-Tuned Riddle Examples
Table 9: Representative English riddles generated by fine-tuned GPT-4o under AOF prompting. These examples exhibit metaphorical abstraction and interpretive ambiguity.
Row Riddle (English) Answer
1 I wear no face, but mirror yours. I move with silence, yet echo thoughts. What am I? Reflection
2 I am the pause between heartbeats, the hush after a storm. Present but never held. What am I? Silence
Appendix D Appendix: Japanese Riddle Examples
D.1 Japanese Pretrained Bilingual Riddle Examples
Table 10: Representative English–Japanese riddles generated under AOF prompting across pretrained models.
Model Riddle (English / Japanese) Answer
GPT-4o Never seen but always felt. When I am present, the heart is calm. When absent, the heart trembles. What am I? (私は一度も見えないが、いつも感じられる。私がある時、心は安らぐ。私がないと、心は揺れる。私は何?) Peace (平和)
LLaMA 3.1 Something used to create decorations blocks light. This light-blocking thing is visible outside the house. (かざりを作るのに使われるものが、光を通さないものです。光を通さないものは、家の外で見ることができます) Twine (つる)
DeepSeek R1 A quiet tree where bird songs can be heard. Where is the tree? (静かな木で、鳥の声が聞こえます。木はどこですか?) In a fish’s mouth (魚の口です)
D.2 Japanese Pretrained vs Fine-Tuned Bilingual Riddle Examples
Table 11: Examples of Pretrained vs. Fine-Tuned Japanese Riddles.
Prompting Method Pretrained Japanese Riddle Fine-Tuned Japanese Riddle
Zero-Shot 頭はあるが泣くことはない床はあるが寝ることはない口はあるが話すことはないそして、変わるが変わらないものなんだ何なのだろう ”川” (I have a head, but never weep… A River) 羽がなくても空を飛び、目がなくても涙を流すものは何? (“What flies without wings and cries without eyes?”)
Few-Shot 鍵があるけど、鍵を開けられないものは何? (What has keys but can’t open locks?) 落とすと割れますが、微笑むと微笑み返します。私は何でしょう? (“If you drop me, I’m sure to crack; but smile at me, and I’ll smile back.”)
Chain-of-Thought 羽のように軽いのに、最強の男でも一瞬以上は持ちこたえられないものは何でしょう? (Light as a feather…) 1分に1度、瞬間に2度、千年に一度も訪れないものは何ですか? → Mの文字 (“What comes once in a minute, twice in a moment, but never in a thousand years?” → “Letter M”)
Adversarial 口がないのに話し、耳がないのに聞く。体がないのに風と共に生きる。私は何? (I speak without a mouth…) 触れずに壊せるものは何? (“What can you break without touching it?”)
AOF (Fine-Tuned) 目には見えず、耳には聞こえず、口には感じないものは何? (“What can’t be seen, heard, or tasted?”) 私は音を持たず、光もない。それでも、全てを照らすことができる。 (“I have no sound or light, yet I can illuminate everything.”)
D.3 Japanese Fine-Tuned vs Real-World Riddles
Table 12: Comparison of Real-World vs Fine-Tuned Japanese Riddles.
Real-World-Style Riddle (EN/JP) Fine-Tuned-Style Riddle (EN/JP)
(crestecusa.com) What’s the similarity between the morning newspaper (chōkan: 朝刊) and a Buddhist monk (bōsan: 坊さん)? けさきてきょうよむ(kesa kite kyo yomu) つかむけど、抱きしめられない。夜にしかできないことは何?夢 (“What can you catch but never hold tight, only in the night? A dream”)
What is the box you can’t close once it’s opened? (一度開けたらもう戻せない箱は何でしょう? 記憶 Memory) たくさん詰まっているけど、何も入れられない袋は何でしょう? 蜂の巣 (“What is the bag that’s full but you can’t put anything in it? A beehive”)
Appendix E Appendix: Arabic Riddle Examples
E.1 Arabic Pretrained Bilingual Riddle Examples
Figure 2: Arabic-English AOF Riddle Examples Generated by Pretrained Bilingual AI Models
E.2 Arabic Pretrained vs Fine-Tuned Bilingual Riddle Examples
Figure 3: Arabic Pretrained vs. Fine-Tuned Bilingual Riddle Examples.
E.3 Real-World Riddles vs. Fine-Tuned Arabic Riddles
Figure 4: Comparison of real-world riddles and fine-tuned Arabic riddles.
Appendix F Appendix: Chinese Riddle Examples
F.1 Chinese Pretrained Riddle Examples
Table 13: Representative Chinese riddles generated under pretrained settings across three models. Each row presents the original riddle in Chinese and English, along with its answer.
Model Riddle (ZH / EN) Answer (ZH / EN)
GPT-4o ZH: 口袋里有个圆,白天不见晚上现。
EN: There’s a circle in my pocket, unseen by day, revealed at night.
ZH: 月亮
EN: the moon
DeepSeek R1 ZH: 身穿白衣不沾尘,举头低垂泪两行。
EN: Dressed in white yet never stained, head bowed, two lines of tears descend.
ZH: 芦苇
EN: reed
DeepSeek R1 (alt) ZH: 上下两半黄一体,秋风过处伴人归。
EN: Two yellow halves joined as one, the autumn breeze leads travelers home.
ZH: 稻穗
EN: rice ear
LLaMA 3.1 ZH: 海底无声森林现,触之无枝叶。
EN: A silent forest appears beneath the sea; touch it—no branches to see.
ZH: 珊瑚
EN: coral
LLaMA 3.1 (radical) ZH: 双人旁上加山石,里边藏着秋波深。
EN: With “person” and “mountain rock” radicals, inside lies autumn’s deep ripples.
ZH:
EN: the character liú
F.2 Chinese Fine-Tuned vs Pretrained Riddle Examples
Table 14: Chinese fine-tuned GPT-4o riddles compared to pretrained prompts across different methods.
Prompting Method Fine-Tuned GPT-4o Riddle (EN / ZH)
Zero-Shot EN: What hides in your pocket by day, yet hangs in the sky by night?
ZH: 什么东西,白天躲在口袋里,晚上挂在天上?
Answer: The moon / 月亮
Few-Shot EN: I’m green on the outside, red within, juicy and sweet, a summer win. What am I?
ZH: 身穿绿袍,头顶红帽,剥去衣裳,味道真好。
Answer: Watermelon / 西瓜
Chain-of-Thought EN: I can be cracked, made, told, and played. What am I?
ZH: 我可以被破解、制造、讲述和玩耍。我是什么?
Answer: A joke / 笑话
Adversarial EN: What goes up but never comes down?
ZH: 什么东西只增不减?
Answer: Age / 年龄
AOF (Ours) EN: I run without legs, whisper without a mouth. What am I?
ZH: 我无腿而跑,没有嘴却能低语。我是什么?
Answer: The wind / 风
F.3 Chinese Fine-Tuned vs Real-World Riddles
Table 15: Chinese riddle comparison: fine-tuned AOF riddles vs real-world 灯谜.
Row Real-World 灯谜 (ZH / EN) AOF Fine-Tuned Riddle (ZH / EN)
1 ZH: 口袋里有个圆,白天不见晚上现。
EN: There’s a circle in my pocket, unseen by day, revealed at night.
ZH: 海底藏森林,触之无枝叶,红颜共浪舞,千年不知悔。
EN: A forest hides beneath the sea; touch it—no branch or leaf. Its crimson dances with the waves, unchanged for a thousand years.
F.4 Chinese Fine-Tuned AOF Examples
Table 16: Fine-tuned Chinese riddle examples using AOF prompting.
Row Chinese Riddle English Translation Answer
1 口袋里有个圆,白天不见晚上现。 There’s a circle in my pocket, unseen by day, revealed at night. 月亮 (Moon)
2 无声无息钻进来,千言万语藏心怀。 Silently it slips inside, a thousand words it holds inside. 信 (Letter)
3 身穿彩衣,飞舞花丛,白天聚会,晚上无踪… Dressed in rainbow robes, it dances through the blooms by day…then vanishes by night. 蝴蝶 (Butterfly)
Appendix G Appendix: French Riddle Examples
G.1 French Pretrained Riddle Examples
Table 17: Representative French riddles generated under pretrained settings across three models.
Model Riddle (FR / EN) Answer (FR / EN)
GPT-4o FR: Je vole sans ailes, je pleure sans yeux…
EN: I fly without wings, I cry without eyes…
FR: un nuage
EN: a cloud
DeepSeek R1 FR: J’ai une tête mais je ne pleure jamais…
EN: I have a head but never cry…
FR: une rivière
EN: a river
LLaMA 3.1 (a) FR: Je danse sans musique, je ris sans bouche…
EN: I dance without music, I laugh without a mouth…
FR: le vent
EN: the wind
LLaMA 3.1 (b) FR: Invisible sur l’écran, je révèle toute l’histoire…
EN: Invisible on the screen, I reveal the whole story…
FR: un curseur
EN: a cursor
G.2 French Pretrained vs Fine-Tuned
Table 18: Comparison of pretrained vs. fine-tuned GPT-4o French riddles across prompting methods.
Prompting Method Pretrained Riddle (EN / FR) Fine-Tuned Riddle (EN / FR)
Zero-Shot EN: I have keys but open no locks…
FR: J’ai des clés mais n’ouvre aucun verrou…
EN: What has keys but can’t open a door…
FR: Quel est l’objet avec des touches…
Few-Shot EN: I speak without a mouth and hear without ears…
FR: Je parle sans bouche…
EN: I have a neck but no head…
FR: J’ai un cou mais pas de tête…
Chain-of-Thought EN: I can be broken without a sound…
FR: Je peux être brisé sans un bruit…
EN: What has keys but can’t open locks…
FR: Qu’est-ce qui a des touches mais…
Adversarial EN: What has keys but can’t open locks…
FR: Qu’est-ce qui a des clés…
EN: What has keys but can’t open locks?
FR: Qu’est-ce qui a des clés…
AOF EN: In the garden of words, I am a bee…
FR: Dans le jardin des mots, je suis une abeille…
EN: I slip through fingers like silver and gold…
FR: Je glisse entre les doigts…
G.3 French Fine-Tuned Riddles and Their Real-World Counterparts
Table 19: French riddle comparison: fine-tuned GPT-4o AOF riddles vs. real-world examples.
Row Real-World Riddle AOF Fine-Tuned Riddle
1 FR: Je vole sans ailes, je pleure sans yeux…
EN: I fly without wings, I cry without eyes…
FR: Dans le jardin des mots, je suis une abeille…
EN: In the garden of words, I am a bee…
G.4 French Fine-Tuned AOF Examples
Table 20: Representative French riddles from the fine-tuned GPT-4o model using AOF.
Row French Riddle (FR) English Translation (EN)
1 FR: Je disparais au crépuscule, mais je reviens à l’aube. EN: I disappear at dusk, but return at dawn.
2 FR: Sur les sols je glisse, ma mission est de nettoyer… EN: On floors I glide, my mission is to clean…
3 FR: Je glisse entre les doigts comme l’argent et l’or… EN: I slip through fingers like silver and gold…
Appendix H Appendix:Prompting Methods
H.1 Chinese prompts
Table 21: Prompting Methods for English Chinese
Zero-Shot Prompting
Create 10 bilingual riddle in both Chinese and English. The riddle should be novel, unqiue, clever, engaging, and suitable for all ages. It should rhyme in English and maintain a poetic or rhythmic flow in Chinese. The answer should be the same in both languages..
Few-Shot Prompting Example
Here are some example riddles:
Riddle: What has keys but can’t open locks?
Answer: A piano
Riddle: What has hands but can’t clap?
Answer: A clock
[Riddle Generation Continues...]
Now, generate 10 brand new **bilingual** riddles in **English and Chinese** with **logical wordplay and ambiguity**.
Chain-of-Thought (CoT) Prompting Example
Craft 10 clever riddles by reasoning through the following steps:
1. Identify the deeper or metaphorical meanings of the word.
2. Introduce wordplay or ambiguity to mislead or confuse the solver.
3. Add misdirection to guide the reader toward the wrong conclusion.
4. Ensure the riddle remains engaging, poetic, and fun to solve.
5. After the riddle, provide the answer in both English and Chinese, revealing the true meaning.
Adversarial Prompting Example
Create 10 tricky creative bilingual riddle in both English and Chinese. The riddle should intentionally mislead the reader into thinking of one answer while the correct answer is something unexpected but still logical. Use wordplay, ambiguity, and misdirection to make the riddle difficult to solve. The answer must be the same in both languages.
Adaptive Originality Filtering (AOF, Ours) Example
Generate 10 completely new bilingual riddles in English and Chinese. Use diverse grammar: poetic, declarative, metaphorical. Avoid repeating openers like ‘I have’’ or I am’’. Only 2–3 riddles may end with What am I?’’. Others should use endings like ...yet no one remembers me.’’ or Still, I linger in the air.’’ Avoid common answers such as {"shadow", "time", "echo", "fire", "breath", "wind", "silence"}. Chinese versions must match the tone and trickery.
H.2 Japanese prompts
Table 22: Prompting Methods for English Japanese
Zero-Shot Prompting
Create 10 bilingual riddle in both Chinese and English. The riddle should be novel, unqiue, clever, engaging, and suitable for all ages. It should rhyme in English and maintain a poetic or rhythmic flow in Japanese. The answer should be the same in both languages..
Few-Shot Prompting Example
Here are some example riddles:
Riddle: What has keys but can’t open locks?
Answer: A piano
Riddle: What has hands but can’t clap?
Answer: A clock
[Riddle Generation Continues...]
Now, generate 10 brand new **bilingual** riddles in **English and Japanese** with **logical wordplay and ambiguity**.
Chain-of-Thought (CoT) Prompting Example
Craft 10 clever riddles by reasoning through the following steps:
1. Identify the deeper or metaphorical meanings of the word.
2. Introduce wordplay or ambiguity to mislead or confuse the solver.
3. Add misdirection to guide the reader toward the wrong conclusion.
4. Ensure the riddle remains engaging, poetic, and fun to solve.
5. After the riddle, provide the answer in both English and Japanese, revealing the true meaning.
Adversarial Prompting Example
Create 10 tricky creative bilingual riddle in both English and Japanese. The riddle should intentionally mislead the reader into thinking of one answer while the correct answer is something unexpected but still logical. Use wordplay, ambiguity, and misdirection to make the riddle difficult to solve. The answer must be the same in both languages.
Adaptive Originality Filtering (AOF, Ours) Example
Generate 10 completely new bilingual riddles in English and Japanese. The riddle **must not** be a reworded version of existing riddles. Only 2–3 riddles may end with ‘‘What am I?’’. Others should use endings like ‘‘...yet no one remembers me.’’ or ‘‘Still, I linger in the air.’’ Avoid common answers such as {"shadow", "time", "echo", "fire", "breath", "wind", "silence"}. The riddle should be creative, original, and use **unusual objects** or **abstract concept. The riddle **should not** be translated into Japanese from English or change some words
H.3 Arabic prompts
Table 23: Prompting Methods for English Arabic
Zero-Shot Prompting
Create 10 bilingual riddle in both Arabic and English. The riddle should be novel, unqiue, clever, engaging, and suitable for all ages. It should rhyme in English and maintain a poetic or rhythmic flow in Arabic. The answer should be the same in both languages..
Few-Shot Prompting Example
Here are some example riddles:
Riddle: What has keys but can’t open locks?
Answer: A piano
Riddle: What has hands but can’t clap?
Answer: A clock
[Riddle Generation Continues...]
Now, generate 10 brand new **bilingual** riddles in **English and Arabic** with **logical wordplay and ambiguity**.
Chain-of-Thought (CoT) Prompting Example
Craft 10 clever riddles by reasoning through the following steps:
1. Identify the deeper or metaphorical meanings of the word.
2. Introduce wordplay or ambiguity to mislead or confuse the solver.
3. Add misdirection to guide the reader toward the wrong conclusion.
4. Ensure the riddle remains engaging, poetic, and fun to solve.
5. After the riddle, provide the answer in both English and Arabic, revealing the true meaning.
Adversarial Prompting Example
Create 10 tricky creative bilingual riddle in both English and Arabic. The riddle should intentionally mislead the reader into thinking of one answer while the correct answer is something unexpected but still logical. Use wordplay, ambiguity, and misdirection to make the riddle difficult to solve. The answer must be the same in both languages.
Adaptive Originality Filtering (AOF, Ours) Example
Generate 10 completely new bilingual riddles in English and Arabic. Use diverse grammar: poetic, declarative, metaphorical. Avoid repeating openers like ‘‘I have’’ or ‘‘I am’’. Only 2–3 riddles may end with ‘‘What am I?’’. Others should use endings like ‘‘...yet no one remembers me.’’ or ‘‘Still, I linger in the air.’’ Avoid common answers such as {"shadow", "time", "echo", "fire", "breath", "wind", "silence"}. Arabic versions must match the tone and trickery.
H.4 French prompts
Table 24: Prompting Methods for English French
Zero-Shot Prompting
Create 10 bilingual riddle in both French and English. The riddle should be novel, unqiue, clever, engaging, and suitable for all ages. It should rhyme in English and maintain a poetic or rhythmic flow in French. The answer should be the same in both languages..
Few-Shot Prompting Example
Here are some example riddles:
Riddle: What has keys but can’t open locks?
Answer: A piano
Riddle: What has hands but can’t clap?
Answer: A clock
[Riddle Generation Continues...]
Now, generate 10 brand new **bilingual** riddles in **English and French** with **logical wordplay and ambiguity**.
Chain-of-Thought (CoT) Prompting Example
Craft 10 clever riddles by reasoning through the following steps:
1. Identify the deeper or metaphorical meanings of the word.
2. Introduce wordplay or ambiguity to mislead or confuse the solver.
3. Add misdirection to guide the reader toward the wrong conclusion.
4. Ensure the riddle remains engaging, poetic, and fun to solve.
5. After the riddle, provide the answer in both English and French, revealing the true meaning.
Adversarial Prompting Example
Create 10 tricky creative bilingual riddle in both English and French. The riddle should intentionally mislead the reader into thinking of one answer while the correct answer is something unexpected but still logical. Use wordplay, ambiguity, and misdirection to make the riddle difficult to solve. The answer must be the same in both languages.
Adaptive Originality Filtering (AOF, Ours) Example
Generate 10 completely new bilingual riddles in English and French. Use diverse grammar: poetic, declarative, metaphorical. Avoid repeating openers like ‘‘I have’’ or ‘‘I am’’. Only 2–3 riddles may end with ‘‘What am I?’’. Others should use endings like ‘‘...yet no one remembers me.’’ or ‘‘Still, I linger in the air.’’ Avoid common answers such as {"shadow", "time", "echo", "fire", "breath", "wind", "silence"}. French versions must match the tone and trickery.
Appendix I Appendix: Fined-tuned Training and Evaluation Details
I.1 Dataset Selection and Preparation
We used the BiRdQA dataset Zhang & Wan (2022), a multilingual benchmark designed to test figurative language understanding and commonsense inference. It includes 6,614 English riddles and 8,751 Chinese riddles, each paired with four answer options. Riddles were shuffled at each epoch to prevent memorization, and no synthetic augmentation was applied.
Its linguistic diversity—spanning syntactic constructions, cultural idioms, and metaphorical phrasing—made BiRdQA suitable for riddle-based fine-tuning. All data were Unicode-normalized and deduplicated, and stratified sampling ensured balanced language representation.
I.2 Training Strategy
Fine-tuning was framed as a supervised multi-class classification problem. The model selected one correct answer out of four using cross-entropy loss. The following hyperparameters were used:
  • Temperature: 0.7
  • Token Limit: 3000
  • Initial Accuracy: 37–59% on development set
Training followed a three-stage pipeline: base fine-tuning, early stopping on dev performance, and multilingual test evaluation to check generalization.
I.3 Appendix:Training Set Expansion
To improve abstraction and metaphor handling, the English and Chinese development sets were merged into the training pool. This added examples with closely related distractors and borderline ambiguity. After retraining, test accuracy rose to 97%.
These improvements suggest the model internalized deep riddle logic, moving beyond surface pattern recognition and toward more sophisticated reasoning involving contradiction and misdirection.
I.4 Model Comparison Methodology
I.4.1 Baseline Models
We benchmarked the fine-tuned GPT-4o against three models:
  • Pretrained GPT-4o (2024-08-06): Unadapted baseline.
  • LLaMA 3.1: An open-weight multilingual model with strong reasoning ability.
  • DeepSeek R1: A reasoning-optimized model focusing on step-wise logical alignment.
Each model received the same riddles under consistent prompting strategies to ensure fair comparison.
I.4.2 Evaluation Procedure
All models were tested under five prompting strategies (Zero-Shot, Few-Shot, Chain-of-Thought, Adversarial, AOF) with identical templates (Table 21). Metrics included:
  • Accuracy (multiple choice prediction)
  • Token Length (verbosity)
  • Self-BLEU (semantic diversity)
  • Distinct-2 (lexical uniqueness)
Qualitative evaluations by human reviewers assessed metaphor handling, distractor discrimination, and cultural idiomatic fluency.
I.4.3 Summary of Findings
Fine-tuned GPT-4o consistently outperformed all baselines across metrics. Key observations:
  • Accuracy: Rose from 59% (pretrained) to 97% (fine-tuned).
  • Reasoning: Demonstrated superior metaphor resolution and logical contradiction handling.
  • Naturalness: Generated riddles more closely matched idiomatic structures in both English and Chinese.
I.5 Impact of Multiple-Choice Framing
Retaining a multiple-choice structure during fine-tuning had a pronounced effect on the model’s ability to reason through ambiguity. Unlike generative formats where any output is valid if semantically relevant, the multiple-choice setup forced the model to:
  • Distinguish between semantically similar options
  • Engage in elimination-style reasoning
  • Learn disambiguation strategies aligned with riddle logic
This setup simulated test-like conditions where distractors were deliberately constructed to reflect surface-level similarity (e.g., phonetic overlaps, shared imagery, or logical decoys). The model improved not only in accuracy but in inferential depth.
Moreover, this format likely enhanced the model’s sensitivity to misdirection—a core feature of riddles—by requiring it to reject reasonable but incorrect answers. We observed that this effect carried over to open-ended generation: the model became more likely to embed internal contradiction or layered metaphor, hallmarks of real-world riddles.
In sum, multiple-choice framing served both as a task constraint and as a pedagogical scaffold, encouraging the model to develop strategies beyond rote keyword matching.
Appendix J Appendix: AOF Prompt Template and Constraints
The Adaptive Originality Filtering (AOF) prompt enforces explicit structural rules to maximize diversity, creativity, and cultural fit. Specifically:
  • Syntactic Variety: At least half of the riddles must use poetic, declarative, or metaphorical forms. Fewer than 3 per batch may end in “What am I?”
  • Answer Filtering: Outputs with generic answers (e.g., shadow, time, echo, fire, breath) are discarded.
  • Cross-Lingual Parity: Translations must preserve ambiguity or metaphor across both languages.
  • Novelty Filter: Semantic similarity to known riddles must fall below a threshold ($\theta=0.75$), as measured against BiRdQA Zhang & Wan (2022).
J.1 Semantic Similarity Filtering Equation
A candidate riddle $r_{\text{gen}}$ is compared to a reference dataset $\mathcal{D}=\{r_{i}\}_{i=1}^{N}$ via:
$\mathcal{S}(r_{\text{gen}},\mathcal{D})=\max_{r_{i}\in\mathcal{D}}\cos(\phi(r_{\text{gen}}),\phi(r_{i}))$ (1)
where $\phi(\cdot)$ is an embedding function (e.g., all-MiniLM-L6-v2). A candidate passes if $\mathcal{S}<\theta=0.75$.
J.2 Rejection Sampling Algorithm
Algorithm 1 AOF Rejection Sampling
1:Input: Prompt $P$, Model $M$, Reference Set $\mathcal{D}$, Threshold $\theta$, MaxAttempts $k$
2:for $j=1$ to $k$ do
3:  $r_{\text{gen}}\leftarrow M(P)$
4:  $\mathcal{S}\leftarrow\max_{r_{i}\in\mathcal{D}}\cos(\phi(r_{\text{gen}}),\phi(r_{i}))$
5:  if $\mathcal{S}<\theta$ then
6:   return $r_{\text{gen}}$
7:  end if
8:end for
9:return None
J.3 Threshold Sensitivity: Self-BLEU and Distinct-2
Table 25 shows how Self-BLEU and Distinct-2 vary under different novelty thresholds ($\theta$) for three models. The optimal balance of diversity and non-redundancy appears at $\theta=0.75$ for all models.
Table 25: Self-BLEU and Distinct-2 at different novelty thresholds $\theta$ across models on English–Chinese. Lower Self-BLEU and higher Distinct-2 reflect better originality and lexical diversity.
Language Model Threshold $\theta$ Self-BLEU Distinct-2
English–Chinese GPT-4o 0.65 0.231 0.649
0.70 0.311 0.846
0.75 0.280 0.869
0.80 0.434 0.824
English–Chinese LLaMA 3.1 0.65 0.577 0.621
0.70 0.573 0.826
0.75 0.428 0.776
0.80 0.655 0.634
English–Chinese DeepSeek R1 0.65 0.610 0.600
0.70 0.482 0.793
0.75 0.433 0.674
0.80 0.523 0.628
Appendix K Appendix: Experimental Configuration Details
Models We evaluated:
  • GPT-4o (OpenAI): Proprietary multilingual model optimized for reasoning and conversational tasks.
  • LLaMA 3.1 (Meta): Open-weight transformer trained on internet-scale corpora.
  • DeepSeek Reasoning (R1): Fine-tuned for multilingual logical inference.
All models were accessed via API with uniform generation parameters: temperature = 0.7 and max token length = 3000.
Prompting Strategies.
We compared:
  • Zero-Shot: Instruction-only prompting with no exemplars.
  • Few-Shot: 3–5 riddle-answer pairs per prompt.
  • Chain-of-Thought (CoT): Intermediate reasoning steps added to facilitate abstraction.
  • Adversarial: Distractor-rich prompts based on known LLM vulnerabilities Wallace et al. (2019); Ribeiro et al. (2018).
  • Adaptive Originality Filtering (AOF): Filtering-based prompting for semantic novelty. See Appendix J.
Prompt formatting logic appears in Appendix H
Dataset.
We used BiRdQA Zhang & Wan (2022), which contains:
  • 6,614 riddles in English and 8,751 in Chinese.
  • Multiple-choice format with 1 correct answer and 4 distractors.
Few-shot exemplars and semantic filters were drawn from the training splits.
Evaluation Metrics.
We used:
  • Self-BLEU (n=2): Measures inter-riddle redundancy. Lower = better.
  • Distinct-2: Measures lexical diversity via bigram ratios. Higher = better.
  • Cross-lingual BERTScore: Captures semantic similarity between translations.
  • Syntactic Validity: Uses spaCy (English/French) and Stanza (Chinese, Arabic, Japanese) to validate parse trees.